4. Assuming a current I in a semi-circular wire with radius a. What is the magnetic flux density on a point at the center ofAssuming a current I in a semi-circular wire with radius a. What is the magnetic flux density on a point at the center of the semi-circle and above the semi-circle plane by a distance d. Also find the magnetic flux density if the circle is full.
Answer #1

-- - Semy Now) dB = a no Idd 41 82 I doi= Mo Idesino we know that 4T 82 do - Mo Iditar i Horia sino = L 41 R12 !! 0 = 90 T ãąDAHULU density (3) at the center of magnetic fleex full wrebe IdBl= Mo Ide sino 4T 82 | Thi sino = 1 0=90 Cuorent element iswhere do = P doão R = 10,0,d) -(x,y,o) k= fapt blåz 1R1 = (p2 +62) de x R = ap ad azi To edo o ttaa 2Ã = pbbdo ar + P2 do âz2 24 Грох – dФа, 41 (4”- е охрана, 2 (440% В - Ког ф2 — и x 4°Ру here eza в юга“ или 244-+aj% CScanned with CamScanner